A force F~ = Fx ˆı + Fy ˆ acts on a particle that
undergoes a displacement of ~s = sx ˆı + sy ˆ
where Fx = 2 N, Fy = −4 N, sx = 1 m, and
sy = 1 m.
Find the work done by the force on the
particle.
Answer in units of J.
Don't round answer
1
Expert's answer
2019-11-04T10:37:59-0500

If the force is constant, the work done by this force is the dot product of force and displacement: "A = \bold F \cdot \bold s". Hence, the work done is "A = (2, -4) \cdot (1, 1) = 2 \cdot 1 + (-4) \cdot 1 = -2 J".